OpenStudy (anonymous):

A traffic light is red for 38 seconds, yellow for 10 seconds, and green for one minute and 12 seconds for cycle. what is the probability that at the moment a passing pedestrian glances at it, the light will be green?

OpenStudy (nurali):

find how many seconds the light is green: 60+12=72. Find the number of seconds it takes to go through one cycle: 38+10+72=120. Divide the seconds it's green by total : 72/120=(3*24)/(5*24)=3/5.